OpenStudy (anonymous):

For the function, find and simplify [f(x + h) − f(x)]/h. (Assume h ≠ 0.) how do I plug in f(x)=9x^2 into [f(x + h) − f(x)]/h

zepdrix (zepdrix):

So we're trying to set up our difference quotient: \[\Large\rm \frac{\color{orangered}{f(x+h)}-\color{royalblue}{f(x)}}{h}\] They provided us with this function:\[\Large\rm \color{royalblue}{f(x)=9x^2}\]Understand the color coding? see where we plug our function in?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\rm \frac{\color{orangered}{f(x+h)}-\color{royalblue}{9x^2}}{h}\]Ok good, that takes care of that part then. Now we need to deal with the \(\Large\rm f(x+h)\)

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

I been confused since I was not given f(x+h)

zepdrix (zepdrix):

So we're taking our function f, and instead of plugging x into it, we're plugging x+h in.\[\Large\rm f(\color{green}{x})=9(\color{green}{x})^2\]In the same way that we might plug 2 in for x,\[\Large\rm f(\color{green}{2})=9(\color{green}{2})^2\]We can also plug in x+h for our x.\[\Large\rm f(\color{green}{x+h})=9(\color{green}{x+h})^2\]Understand how that works? :d

OpenStudy (anonymous):

Why is it that for f(x+h) we plug (x+h) into 9x^2 but for f(x) we left it as 9x^2.

OpenStudy (anonymous):

I am confused on how f(x+h) is 9(x+h)^2

zepdrix (zepdrix):

Function notation can be a little tricky :) here is our function:\[\Large\rm f(~~)=9(~~)^2\]Just think of it like this for a moment.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

So when we say that this is a function of x, we're "plugging" x into our function,\[\Large\rm f(\color{orangered}{x})=9(\color{orangered}{x})^2\]When we plug something different in,\[\Large\rm f(\color{orangered}{x+h})\]We get something different out,\[\Large\rm =9(\color{orangered}{x+h})^2\]

zepdrix (zepdrix):

When it was written as a function of x, the x's are x. :P When we instead plug in x+h, we x's are replaced with x+h's. Still a lil confusing? :d

OpenStudy (anonymous):

oh wow ok I like that. Makes sense

zepdrix (zepdrix):

So here is our function `evaluated at` (value plugged in) x+h, \[\Large\rm \color{orangered}{f(x+h)=9(x+h)^2}\]

OpenStudy (anonymous):

so here is what i get when I actually plug f(x) and f(x+h) into the equation without simplifying: [9(x+h)^2 - 9x^2] / h

zepdrix (zepdrix):

Ok great!

zepdrix (zepdrix):

Gotta be careful with your simplification. Remember how to expand a binomial?

OpenStudy (anonymous):

and when I have simplified I came up with this: [9(x^2+2xh+h^2)-9x^2] / h then : (9x^2+18xh+9h^2-9x^2) / h

OpenStudy (anonymous):

then : (18xh+9h^2) / h

OpenStudy (anonymous):

can I then factor out 9h and be left with : [9h(2xh+h)] h ?

OpenStudy (anonymous):

oh I cant factor out 9h can I?

zepdrix (zepdrix):

Yes you can. But there is one further simplification step from there. Do the numerator and denominator share any factors? Can we do anything about that?

OpenStudy (anonymous):

yes, H

OpenStudy (anonymous):

ok : so..... if I did that before pulling out 9h I would have 18x+9h ?

zepdrix (zepdrix):

Yessss, very good \c:/

zepdrix (zepdrix):

You can pull out the 9 from there if you want, but it's not necessary.

OpenStudy (anonymous):

ok I have one mroe quick question and I will take it from here if I am correct ok

OpenStudy (anonymous):

from what you have taught me I will attempt to plug in the next two functions on my own

OpenStudy (anonymous):

I have f(x) = 7x2 − 9x + 3 so [f(x + h) − f(x)] / h so I would have {[7(x+h)^2-9(x+h)+3]-[7x^2-9x+3]} / h

OpenStudy (anonymous):

?

OpenStudy (anonymous):

f(x) 7x2 − 9x + 3

zepdrix (zepdrix):

Yessss, very good on your setup. Remember that in order to drop these brackets [7x^2-9x+3], you'll have to distribute the negative to each term.

OpenStudy (anonymous):

yes I did and I simplified and I ended up with this (7x^2+14xh+9h+9x+7h^2) / h

OpenStudy (anonymous):

then if i simplified it with H being a common denominator I came up with this: 7x^2+14x+9+9x+7h 7x^2+23x+7h+9 , is this correct?

zepdrix (zepdrix):

Before you do our division, taking the h's out, You shouldn't be left with any non-h terms. Just something to keep in mind. So I think we made a little mistake somewhere. Let's see where.

zepdrix (zepdrix):

So from here, \[\Large\rm \frac{7(x+h)^2-9(x+h)+3-7x^2+9x-3}{h}\]Expanding gives us:\[\Large\rm \frac{7x^2+14xh+h^2-9x-9h+3-7x^2+9x-3}{h}\]Combining like-terms:\[\Large\rm \frac{\cancel{7x^2}+14xh+h^2\cancel{-9x}-9h+\cancel{3}\cancel{-7x^2}+\cancel{9x}\cancel{-3}}{h}\]

zepdrix (zepdrix):

Early on you had a -9(x+h) and I think you accidentally turned it into 9x+9h. It should have become -9x-9h

OpenStudy (anonymous):

oh wow, I should have ruled out the common denominator beforehand like you are showing, that would save a lot of work. and yes I caught that thank you:)

OpenStudy (anonymous):

well after erasing opposing "factors" ez: -3 , +3

OpenStudy (anonymous):

ok... I see what you did

OpenStudy (anonymous):

I can simplify h from the numerator, and it ends up canceling out with the H from the denominator ?right?

OpenStudy (anonymous):

leaving me with 14xh +h -9

zepdrix (zepdrix):

Yes, everyone in the top loses an h. Woops I think you forgot to take an h out of 14xh

OpenStudy (anonymous):

yes you're right, this is why its bad to work and do my homework at the same time. hah

zepdrix (zepdrix):

haha XD

OpenStudy (anonymous):

at least I have the main issue down , how to plug f(x) and f(h+x) into an equation, now I just have to go slowly through the steps and always go back and recheck those steps. :)

OpenStudy (anonymous):

So I plugged in our final answer, and was told it was wrong.

zepdrix (zepdrix):

good good good :) your basic algebra steps seem very good. I'm glad we got past that function stuff without too much trouble.

zepdrix (zepdrix):

Wrong? +_+ Lies... Sec I'll rework it on paper :3

OpenStudy (anonymous):

me too even though you will prbably have it done way before I do

zepdrix (zepdrix):

This is what the function looks like?\[\Large\rm f(x)= 7x^2 − 9x + 3\]

zepdrix (zepdrix):

\[\Large\rm f(x)= 7x^2 -9x + 3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and f(x+h) - f(x) all over h

zepdrix (zepdrix):

14x-9+7h That's what I'm coming up with ^ Ahhh I think I didn't look at your answer closely enough last time >.<

zepdrix (zepdrix):

Oh oh oh that was my fault. I didn't distribute the 7 to the h^2. that term in the numerator should had been a 7h^2.

OpenStudy (anonymous):

I eneded up with this: 14xh+7h^2-9h

zepdrix (zepdrix):

Before taking an h out? Good.

OpenStudy (anonymous):

hahaha i didnt see your psot, but I missed it too so its ok, its a long equation and its late a night,. its forgivable by me lol

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then : 14x+7h-9

OpenStudy (anonymous):

and I don't believe I can simplify that anymore, so This will be my final answer

OpenStudy (anonymous):

yes!!!! now it is correct!! thank you@zepdrix for all your help! :)

zepdrix (zepdrix):

Yay team \c:/